# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

### v2

@@ IV-B Implementation
Base Model. In ToolGen, we explore the utilization of three distinct code LLMs to encompass diverse model architectures and parameter scales. These code LLMs demonstrate impressive performance in code generation and have found extensive utilization in prior studies [23, Compilable, 4, 5, 12] for fine-tuning and evaluation.
@@
Trigger Insertion. To create the augmented dataset for fine-tuning the employed base model, we begin with the Python corpus from the CodeSearchNet dataset. Since the CodeSearchNet dataset does not provide complete code repositories from which to extract Python functions, we initiate the process by crawling the code repositories listed in the dataset. Subsequently, we follow the procedure outlined in Section III-B1 to extract and augment functions within these code repositories, ultimately generating the augmented dataset. It’s important to note that the CodeSearchNet dataset includes a partitioning into training, validation, and test sets. For our trigger insertion process, we exclusively utilize the code repositories associated with the training set. The resulting augmented dataset comprises a total of 256,870 pairs of docstrings and augmented functions, which are sourced from 12,231 distinct Python code repositories. Regarding dataset statistics, the average token count for docstrings is 10.96, and for augmented functions, it is 54.13. Additionally, the special token <COMP> appears an average of 5.37 times within these functions.
Model Fine-tuning. In the fine-tuning process, we adopt different strategies for CodeGPT, CodeT5, and CodeLlama: For CodeGPT and CodeT5, we perform full-parameter fine-tuning, optimizing all model parameters during this phase. In the case of CodeLlama, we employ LoRA with a reduction factor (r) of 16 and a scaling factor (alpha) of 16 to achieve parameter-efficient fine-tuning. This approach allows us to optimize only 0.25% of the trainable parameters in comparison to the original CodeLlama model. The fine-tuning settings for learning rate and batch size are consistent across all three models, with a learning rate of 5E-6 and a batch size of 32. However, the number of epochs differs: 10 epochs for CodeGPT and CodeT5, while CodeLlama undergoes fine-tuning for 2 epochs. To ensure reproducibility, we set the seed for random functions to 42 consistently across all packages and libraries used.
@@ IV-C Evaluation Benchmark
IV-C1 Dataset
The construction of our benchmark dataset follows this process: We begin by crawling the code repositories listed in the test set of the CodeSearchNet dataset. Next, we extract pairs of docstrings and functions from these repositories by parsing and traversing Abstract Syntax Trees (ASTs), similar to the method described in Section III-B1. This process yields an evaluation dataset comprising 12,744 Python functions sourced from 680 code repositories. On average, the docstrings contain 10.63 tokens, while the functions consist of an average of 54.56 tokens.
@@ IV-C2 Baselines
We develop three baseline approaches by fine-tuning these same base models but performing straightforward code generation without tool integration. The different variants of ToolGen and the corresponding baselines are presented in Table I, along with the base models they employ.
Specifically, the fine-tuning process for the baselines involves using the 256,870 pairs of docstrings and functions from the augmented dataset. Notably, the fine-tuning is conducted on the original functions, prior to the introduction of <COMP> tokens. The training configurations, including learning rates and training epochs, mirror those employed in the implementation of ToolGen. After fine-tuning, these models are utilized to perform straightforward code generation, as outlined in Section II-A. We label the three baseline approaches as follows:
  • Base-gpt: Represents straightforward code generation using the CodeGPT model fine-tuned on original functions.
  • Base-t5: Signifies straightforward code generation using the CodeT5 model fine-tuned on original functions.
  • Base-llama: Designates straightforward code generation with the CodeLlama model fine-tuned on original functions.
TABLE I: Variants of ToolGen and corresponding baselines.
  Approach Base Model
  Base-gpt CodeGPT
@@
Base-t5 CodeT5
@@
Base-llama CodeLlama
@@ IV-C3 Metrics
In our evaluation, we employ a combination of commonly used similarity metrics and introduce two novel metrics tailored specifically to repository-level code generation.
Similarity Metrics: We utilize the following well-established similarity metrics to measure the correspondence between generated functions and their ground-truth counterparts:
@@
  • Edit Similarity (EditSim) [29]: This metric measures the similarity between two pieces of functions by analyzing the edit operations required to transform one into the other. It provides insights into code structure alignment.
Repository-level Metrics: To assess the effectiveness of both ToolGen and the baselines in repository-level code generation, we introduce the Repository-level metrics, namely Dependency Coverage (DepCov) and Error Rate (ErrRate).
  • Dependency Coverage (DepCov): This metric calculates the ratio of repository-level dependencies, including user-defined functions and attributes, that appear in ground-truth functions and are covered by the generated functions. The process involves static analysis using Jedi to identify expressions (such as function calls and attribute accesses) that can be traced back to their definitions within the code repository. We collect these expressions as a set and evaluate the extent to which they are addressed by the corresponding generated function. For example, if the attribute definition for “self._registered_updates” can be inferred for a ground-truth function, we examine whether it is properly addressed in the corresponding generated function.
  • Static Validity Rate (ValRate): As repository-level dependencies can potentially introduce dependency errors in generated functions, we introduce the Static Validity Rate metric (ValRate) to evaluate the effectiveness of ToolGen in reducing dependency errors. This metric evaluates the proportion of generated functions that successfully pass a static check for dependency errors, specifically no-member and undefined-variable. To perform this evaluation, we incorporate the generated functions into their respective code repositories and conduct static lint analysis using pylint [30]. Functions that do not exhibit syntax errors, no-member, or undefined-variable errors are deemed statically valid. It’s worth noting that this work does not address logic errors, as their identification typically necessitates the execution of corresponding unit tests.
@@ V-A RQ1: Similarity Metrics
The evaluation results of similarity metrics are presented in Table II. When comparing the performance of ToolGen’s variants and the three different base models, namely CodeGPT, CodeT5, and CodeLlama, we observe that ToolGen achieves similarity scores comparable to the baselines.
@@
TABLE II: Evaluation results for similarity metrics. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Similairty Metrics
BLEU-4
CodeBLEU
EditSim
  Base-gpt
0.304
0.284
47.6%
ToolGen-gpt (ours)
0.315 ($\Delta=+3.6\%$)
0.275 ($\Delta=-3.2\%$)
47.1% ($\Delta=-1.1\%$)
Base-t5
0.355
0.291
63.7%
ToolGen-t5 (ours)
0.362 ($\Delta=+2.0\%$)
0.288 ($\Delta=-1.0\%$)
62.3% ($\Delta=-2.2\%$)
Base-llama
0.368
0.310
65.7%
ToolGen-llama (ours)
0.371 ($\Delta=+0.8\%$)
0.295 ($\Delta=-4.8\%$)
62.4% ($\Delta=-5.0\%$)
@@
To provide a detailed breakdown, when utilizing CodeGPT as the base model, ToolGen-gpt demonstrates a 3.6% improvement in BLEU-4 compared to Base-gpt. However, it exhibits a 3.2% decrease in CodeBLEU and a 1.1% decrease in EditSim. Similarly, with the base model CodeT5, ToolGen-t5 exhibits a 2.0% enhancement in BLEU-4 relative to Base-t5 but experiences a 1.0% reduction in CodeBLEU and a 2.2% decrease in EditSim. In the case of the larger base model CodeLlama, ToolGen-llama shows an improvement of 0.8% in BLEU-4 compared to Base-t5 but encounters a 4.8% decrease in CodeBLEU and a 5.0% decrease in EditSim.
@@
SUMMARY: ToolGen demonstrates competitive performance in similarity metrics with the baselines across various base models.
V-B RQ2.a: Dependency Coverage
@@
Table III displays the evaluation results for repository-level Dependency Coverage (DepCov). Notably, our approach ToolGen demonstrates significant superiority over the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen surpasses the corresponding baselines in Dependency Coverage by 45.8%, 37.3%, and 15.2%, respectively.
@@
These results underscore the effectiveness of the tool-integrated generation process in enhancing awareness of repository-level dependencies, a challenge often unaddressed by the conventional code LLMs. For instance, consider the incomplete function in Figure 1: in a straightforward CodeLlama generation, it fails to recognize the valid attributes of “self”. However, through tool-integrated generation, ToolGen leverages Jedi to deduce a list of completion suggestions, enabling it to select the most suitable one and cover target repository-level dependencies, including the usage of user-defined functions and attributes.
@@
SUMMARY: Our approach, ToolGen, consistently outperforms the baselines in repository-level Dependency Coverage across all three base models by ranging from 15.2% to 45.8%. These results highlight the effectiveness of our tool-integrated generation process in addressing the crucial issue of enhancing awareness of repository-level dependencies, which is often a challenge for conventional code LLMs in repository-level code generation.
TABLE III: Evaluation results for repository-level metrics. DepCov and ValRate represent Dependency Coverage and Static Validity Rate, respectively. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Repository-level Metrics
DepCov
ValRate
  Base-gpt
8.3%
32.2%
ToolGen-gpt (ours)
12.1% ($\Delta=+45.8\%$)
35.7% ($\Delta=+10.9\%$)
Base-t5
11.0%
48.6%
ToolGen-t5 (ours)
15.1% ($\Delta=+37.3\%$)
69.1% ($\Delta=+42.2\%$)
Base-llama
13.8%
47.8%
ToolGen-llama (ours)
15.9% ($\Delta=+15.2\%$)
67.1% ($\Delta=+40.4\%$)
V-C RQ2.b: Static Validity Rate
Table III also presents the evaluation results for Static Validity Rate (ValRate) in repository-level lint analysis, with a particular focus on no-member and undefined-variable errors. Remarkably, our approach, ToolGen, consistently exhibits significantly higher Static Validity Rate compared to the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen increases the Static Validity Rate by 10.9%, 42.2%, and 40.4%, respectively.

### v3

@@ IV-B Implementation
Base Model. In ToolGen, we explore the utilization of three distinct code LLMs to encompass diverse model architectures and parameter scales. These code LLMs demonstrate impressive performance in code generation and have found extensive utilization in prior studies [33, 39, 4, 5, 12] for fine-tuning and evaluation.
@@
Trigger Insertion. To create the augmented dataset for fine-tuning the employed base model, we begin with the Python corpus from the training set of CodeSearchNet dataset. Since the CodeSearchNet dataset does not provide complete code repositories from which to extract Python functions, we initiate the process by crawling the code repositories listed in the dataset. Subsequently, we follow the procedure outlined in Section III-B1 to extract and augment functions within these code repositories, ultimately generating the augmented dataset. It’s important to note that the CodeSearchNet dataset includes a partitioning into training, validation, and test sets. For our trigger insertion process, we exclusively utilize the code repositories associated with the training set. The resulting augmented dataset comprises a total of 249,298 pairs of descriptions and augmented functions, which are sourced from 12,231 distinct Python code repositories. Regarding dataset statistics, the average token count for descriptions is 10.98, and for augmented functions, it is 55.31. Additionally, the special token <COMP> appears an average of 5.54 times within these functions.
Model Fine-tuning. In the fine-tuning process, we adopt different strategies for CodeGPT, CodeT5, and CodeLlama: For CodeGPT and CodeT5, we perform full-parameter fine-tuning, optimizing all model parameters during this phase. In the case of CodeLlama, we employ LoRA with a reduction factor (r) of 8 and a scaling factor (alpha) of 16 to achieve parameter-efficient fine-tuning. This approach allows us to optimize only 3.86% of the trainable parameters in comparison to the original CodeLlama model. The fine-tuning settings for learning rate and batch size are consistent across all three models, with a learning rate of 5E-6 and a batch size of 32. However, the number of epochs differs: 10 epochs for CodeGPT and CodeT5, while CodeLlama undergoes fine-tuning for 3 epochs. To ensure reproducibility, we set the seed for random functions to 42 consistently across all packages and libraries used.
@@ IV-C Evaluation Benchmark
IV-C1 Datasets
To evaluate ToolGen, we curate two datasets: (i) a large dataset derived from the CodeSearchNet [40] to assess similarity-based and dependency-based effectiveness (RQ1 and RQ2); (ii) a dataset derived from CoderEval [27] containing test cases to evaluate execution-based effectiveness (RQ3).
CodeSearchNet. To assess similarity-based and dependency-based effectiveness, we follow this process to construct the dataset: We start by crawling the code repositories listed in the test set of the CodeSearchNet dataset, ensuring no overlap with the training set used for model fine-tuning. We then extract pairs of descriptions and functions from these repositories by parsing and traversing Abstract Syntax Trees (ASTs), similar to the method described in Section III-B1. This process yields an evaluation dataset comprising 12,406 Python functions sourced from 671 code repositories. On average, the descriptions contain 10.66 tokens, while the functions consist of an average of 54.54 tokens.
CoderEval. To evaluate execution-based effectiveness, we initially gather all 230 Python code generation tasks from the CoderEval benchmark, extracted from 43 real-world Python repositories. Each task consists of a natural language description, a ground-truth code snippet, and a set of test cases, along with the project environment context associated with the task (e.g., project source code, dependent libraries, and test scripts). The tasks are categorized into six runnable levels: self-contained, slib-runnable, plib-runnable, class-runnable, file-runnable, and project-runnable [27]. Each runnable level relies on the dependencies defined at that level and does not depend on those defined at subsequent levels. For example, plib-runnable indicate that the task requires public third-party libraries, while file-runnable require dependencies defined in the current file (e.g., user-defined classes and functions). We remove the tasks overlapping with the training dataset of ToolGen, resulting in a final dataset containing 176 tasks.
@@ IV-C2 Baselines
The different variants of ToolGen and the baselines are presented in Table I, along with the base models they employ.
Vanilla Baselines. We develop three vanilla baseline approaches by fine-tuning these same base models but performing straightforward code generation without tool integration. Specifically, the fine-tuning process for the baselines involves using the 249,298 pairs of descriptions and functions from the augmented dataset. Notably, the fine-tuning is conducted on the original functions, prior to the introduction of <COMP> tokens. The training configurations, including learning rates and training epochs, mirror those employed in the implementation of ToolGen. After fine-tuning, these models are utilized to perform straightforward code generation, as outlined in Section II-A. We label the three baseline approaches as follows:
  • Vanilla-gpt: Represents straightforward code generation using the CodeGPT model fine-tuned on original functions.
  • Vanilla-t5: Signifies straightforward code generation using the CodeT5 model fine-tuned on original functions.
  • Vanilla-llama: Designates straightforward code generation with the CodeLlama model fine-tuned on original functions.
Retrieval-Augmented-Generation (RAG) Baselines. We also include RepoCoder [41], a state-of-the-art approach that addresses repository-level code generation by integrating a similarity-based retriever and a pre-trained code language model in an iterative retrieval-augmented-generation pipeline. Similarly, we create three variants of RepoCoder based on the three fine-tuned models Vanilla-gpt, Vanilla-t5, and Vanilla-llama. We directly apply the prompt template defined in the original implementation of RepoCoder. The three variants of RepoCoder are listed as follows:
  • RepoCoder-gpt: Represents the variant of RepoCoder with the CodeGPT model fine-tuned on original functions.
  • RepoCoder-t5: Signifies the variant of RepoCoder with the CodeT5 model fine-tuned on original functions.
  • RepoCoder-llama: Designates the variant of RepoCoder with the CodeLlama model fine-tuned on original functions.
The hyperparameters of RepoCoder used in our experiments follow its default implementation: the retrieval-generation iteration is set to 2, the window size is 20, and the slice size is 2.
RAG-based Variants of ToolGen. In fact, RAG method is orthogonal to our tool-integrated approach. To ensure a fair comparison and further explore the potential of ToolGen, we also develop three RAG-based variants: ragToolGen-gpt, ragToolGen-t5, and ragToolGen-llama. In these variants, the employed retrieval process is the same as the RepoCoder baselines.
TABLE I: Variants of ToolGen and baselines.
  Approach Base Model Architecture # Parameters
  Vanilla-gpt CodeGPT Decoder-Only 124 Million
RepoCoder-gpt
@@
ragToolGen-gpt (ours)
Vanilla-t5 CodeT5 Encoder-Decoder 220 Million
RepoCoder-t5
@@
ragToolGen-t5 (ours)
Vanilla-llama CodeLlama Decoder-Only 7 Billion
RepoCoder-llama
@@ IV-C3 Metrics
In our evaluation, we employ commonly used similarity-based metrics, two novel dependency-based metrics, and an execution-based metric to evaluate the effectiveness of ToolGen in repository-level code generation.
Similarity-based Metrics: We utilize the following well-established similarity metrics to measure the correspondence between generated functions and their ground-truth counterparts:
@@
  • Edit Similarity (EditSim) [43]: This metric measures the similarity between two pieces of functions by analyzing the character-level edit operations required to transform one into the other.
  • Exact Match: This metric measures the ratio of the generated code that are exactly matched with the ground truth.
The calculation of the similarity-based metrics follows the implementation in CodeXGLUE [^4] https://github.com/microsoft/CodeXGLUE.
Dependency-based Metrics: To assess the effectiveness of both ToolGen and the baselines in repository-level code generation, we introduce the dependency-based metrics, namely Dependency Coverage (DepCov) and Static Validity Rate (ErrRate).
  • Dependency Coverage (DepCov): This metric calculates the ratio of repository-level dependencies, including user-defined functions and attributes, that appear in ground-truth functions and are covered by the generated functions. Given the $i$-th ground-truth function $gt_{i}$, we identify dependencies by performing the Trigger Insertion procedure (Algorithm 1) and extracting expressions (such as function calls and attribute accesses like “self._registered_updates”) that are marked with a trigger. These expressions are considered dependencies as their definitions can be traced in the current repository using static analysis tools like Jedi. Next, for the generated function $pred_{i}$ corresponding to $gt_{i}$, we extract all expressions by traversing its corresponding AST. We denote the identified dependencies in $gt_{i}$ and extracted expressions in $pred_{i}$ as two sets, $DEP(gt_{i})$ and $EXP(pred_{i})$, respectively. The Dependency Coverage can be calculated as follows, where $N$ is the size of the test dataset:
    $DepCov=\frac{\sum_{i}^{N}|EXP(pred_{i})\cap DEP(gt_{i})|}{\sum_{i}^{N}|DEP(gt_{i})|}$
  • Static Validity Rate (ValRate): As repository-level dependencies can potentially introduce dependency errors in generated functions, we introduce the Static Validity Rate metric (ValRate) to evaluate the effectiveness of ToolGen in reducing dependency errors. This metric evaluates the proportion of generated functions that successfully pass a static check for dependency errors, specifically no-member and undefined-variable. To perform this evaluation, we incorporate the generated functions into their respective code repositories and conduct static lint analysis using pylint [44]. Functions that do not exhibit syntax errors, no-member, or undefined-variable errors are deemed statically valid. The Static Validity Rate can be calculated as follows:
    $ValRate=\frac{|\{pred_{i\leq N}:pred_{i}~{}\text{passes lint check}\}|}{N}$
Execution-based Metric: To further assess the functional correctness of the generated functions, we also employ a widely used execution-based metric involving running test cases.
  • Test Pass Rate (Pass@1): This metric calculates the ratio of generated functionally-correct functions that pass all corresponding test cases. It is evaluated specifically on the CoderEval dataset, where test cases and test scripts are provided.
@@ V-A RQ1: Similarity-based Effectiveness
The evaluation results of similarity-based metrics are presented in Table II. When comparing the performance of ToolGen’s variants and the three different base models, namely CodeGPT, CodeT5, and CodeLlama, we observe that ToolGen achieves similarity scores comparable to the baselines.
@@
TABLE II: Evaluation results of similarity-based effectiveness. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Similarity-based Metrics
BLEU-4 CodeBLEU EditSim ExactMatch
  Vanilla-gpt 0.331 0.313 65.4% 4.2%
ToolGen-gpt 0.340 ($\Delta=+2.7\%$) 0.310 ($\Delta=-1.0\%$) 64.7% ($\Delta=-1.1\%$) 4.7% ($\Delta=+11.9\%$)
Vanilla-t5 0.341 0.289 63.9% 4.3%
ToolGen-t5 0.362 ($\Delta=+6.2\%$) 0.293 ($\Delta=+1.4\%$) 61.9% ($\Delta=-3.1\%$) 5.5% ($\Delta=+27.9\%$)
Vanilla-llama 0.408 0.360 67.9% 5.7%
ToolGen-llama 0.425 ($\Delta=+4.2\%$) 0.358 ($\Delta=-0.6\%$) 66.3% ($\Delta=-2.4\%$) 6.9% ($\Delta=+21.1\%$)
@@
To provide a detailed breakdown, when utilizing CodeGPT as the base model, ToolGen-gpt demonstrates a 2.7% improvement in BLEU-4 and a 11.9% improvement in Exact Match compared to Vanilla-gpt. However, it exhibits a 1.0% decrease in CodeBLEU and a 1.1% decrease in Edit Similarity. With the base model CodeT5, ToolGen-t5 exhibits 6.2%, 1.4%, and 27.9% enhancements in BLEU-4, CodeBLEU, and Exact Match relative to Vanilla-t5 but experiences a 3.1% decrease in Edit Similarity. In the case of the larger base model CodeLlama, ToolGen-llama shows improvements of 4.2% and 21.1% in BLEU-4 and Exact Match compared to Vanilla-t5 but encounters a 0.6% decrease in CodeBLEU and a 2.4% decrease in Edit Similarity.
Although the absolute improvements in Exact Match rate are not large (from 0.5% to 1.2%), considering the size of the test set (e.g., 12,406 samples), the additional exactly matched functions range from 62 to 149. The variability in ToolGen’s performance across BLEU-4 and CodeBLEU can be attributed to the tokenization methods used in the calculations. For BLEU-4, before using the widely used utility script  [^5] https://github.com/microsoft/CodeXGLUE/blob/main/Text-Code/text-to-code/evaluator/bleu.py, we first tokenize the generated function and its corresponding ground-truth function using the tokenizer of the base code LLMs. In contrast, CodeBLEU is calculated based on the original generated and ground-truth code using the utility script [^6] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/code-to-code-trans/evaluator/CodeBLEU that employs a simpler method, splitting functions into strings (e.g.,func(arg1,)”) based on spaces. This splitting method may introduce inaccuracies in the statistics of matched n-grams, consequently affecting the CodeBLEU calculation. For Edit Similarity, it is calculated at the character level, making it overly sensitive to semantics-insensitive elements like temporary variables. When two variables have different names, their similarity is much lower at the character level than at the token level.
@@
SUMMARY: ToolGen demonstrates competitive performance in similarity metrics compared to the baselines across various base models. It achieves improvements in BLEU-4 and Exact Match while exhibiting comparable performance in CodeBLEU and Edit Similarity.
V-B RQ2: Dependency-based Effectiveness
V-B1 Dependency Coverage
@@
Table III displays the evaluation results for repository-level Dependency Coverage (DepCov). Notably, our approach ToolGen demonstrates significant superiority over the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen surpasses the corresponding baselines in Dependency Coverage by 39.1%, 36.4%, and 31.4%, respectively.
TABLE III: Evaluation results of dependency-based effectiveness. DepCov and ValRate represent Dependency Coverage and Static Validity Rate, respectively. ValRate-dep represents the Static Validity Rate rate calculated only for functions containing dependencies. $\Delta$ indicates the metric improvement or reduction of ToolGen’s variants compared to the baselines.
  Approach Dependency-based Metrics
DepCov ValRate ValRate-dep
  Vanilla-gpt 8.7% 50.4% 46.5%
ToolGen-gpt 12.1% ($\Delta=+39.1\%$) 79.5% ($\Delta=+57.7\%$) 78.0% ($\Delta=+67.7\%$)
Vanilla-t5 11.0% 47.3% 42.5%
ToolGen-t5 15.0% ($\Delta=+36.4\%$) 70.6% ($\Delta=+49.3\%$) 68.0% ($\Delta=+60.0\%$)
Vanilla-llama 14.0% 49.7% 44.4%
ToolGen-llama 18.4% ($\Delta=+31.4\%$) 72.0% ($\Delta=+44.9\%$) 69.6% ($\Delta=+56.8\%$)
@@
These results underscore the effectiveness of the tool-integrated generation process in enhancing awareness of repository-level dependencies, a challenge often unaddressed by the conventional code LLMs. For instance, consider the incomplete function in Figure 1: in a straightforward CodeLlama generation, it fails to recognize the valid attributes of “self”. However, through tool-integrated generation, ToolGen leverages Jedi to deduce a list of completion suggestions, enabling it to select the most suitable one and cover target repository-level dependencies, including the usage of user-defined attributes.
@@
SUMMARY: Our approach, ToolGen, consistently outperforms the baselines in repository-level Dependency Coverage across all three base models by ranging from 31.4% to 39.1%. These results highlight the effectiveness of our tool-integrated generation process in addressing the crucial issue of enhancing awareness of repository-level dependencies, which is often a challenge for conventional code LLMs in repository-level code generation.
V-B2 Static Validity Rate
Table III also presents the evaluation results for Static Validity Rate (ValRate and ValRate-dep) in repository-level lint analysis, with a particular focus on no-member and undefined-variable errors. Remarkably, our approach, ToolGen, consistently exhibits significantly higher Static Validity Rate compared to the baselines across all three base models. Specifically, when employing the base models CodeGPT, CodeT5, and CodeLlama, ToolGen increases the Static Validity Rate (ValRate) by 57.7%, 49.3%, and 44.9%, respectively. When considering only the functions containing dependencies, ToolGen improves the Static Validity Rate (ValRate-dep) by 67.7%, 60.0%, and 56.8%, respectively.